OpenStudy (agent47):

@satellite73, last one. Find E(X|Y=y)

OpenStudy (agent47):

Let (X,Y) have joint mass function as seen in the attachment. Find E(X|Y=y)

OpenStudy (anonymous):

now you are making me think

OpenStudy (agent47):

Yes, this is the last remaining problem in my set :/

OpenStudy (anonymous):

\[E(X|Y=y)=\sum_xxF_{x|y}(x|y)\] if i recall

OpenStudy (anonymous):

unless you have a different formula to use

OpenStudy (agent47):

no that's correct. I've been doing that here as well.. http://openstudy.com/users/agent47#/updates/56c9602ee4b09397a170926f

OpenStudy (anonymous):

link is not working

OpenStudy (anonymous):

you got an example?

OpenStudy (agent47):

Does this help?

OpenStudy (agent47):

That was for this question

OpenStudy (anonymous):

yeah but i have no idea how to apply it in your case here

OpenStudy (agent47):

I think we need to find the general formula for P(X|Y=n) for n = 1, 2, .. k What I've been doing so far was just plugging in numbers and trying to catch a pattern

OpenStudy (anonymous):

the only thing is see in your example is that the final answer should be \[\frac{y+1}{2}\]

OpenStudy (anonymous):

i mean the example you already did, not this one here

OpenStudy (agent47):

Hang on, sorry I may have missed something, and yes that's what i simplified my earlier example to.

OpenStudy (agent47):

That's the whole example, the question is asking me to compute E(X|Y=y) for that example.

OpenStudy (anonymous):

yikes can you show me what \[E(X|Y=1)\] is?

OpenStudy (agent47):

sec, let me try computing it..

OpenStudy (anonymous):

i guess what i am asking is, do we replace n by 1 and sum over k , or what?

OpenStudy (agent47):

Isn't\[P(X=k|Y=n)\] just 0 when n > k and what they have when n <=k ?

OpenStudy (anonymous):

my problem is that to be honest i don't know how to read this

OpenStudy (anonymous):

To \(k,n\) correspond to \(X,Y\)?

OpenStudy (agent47):

So given a joint pmf, find E(X|Y=y).. to make this somewhat simpler for myself I can just say E(X=k|Y=n) maybe?\[P(X=k|Y=n)=C*2^{-k}/n, \space n \le k \space and \space 0 \space otherwise\] @wio i think so

OpenStudy (anonymous):

And is \(E\) expected value?

OpenStudy (agent47):

\[E(X=k|Y=n)=\sum_{k=1}^{\infty}k*P(X=k|Y=n)\]@wio yes

OpenStudy (anonymous):

just dug out my copy of Ross and that is exactly what it says

OpenStudy (agent47):

Yea it's a screenshot from A Natural intro to probability

OpenStudy (anonymous):

so now my only question is what is \[P(X=k|Y=n)\] in this case?

OpenStudy (agent47):

\[P(X=k \cap Y=n)/P(Y=n)\]

OpenStudy (agent47):

I think the top is already given, we just need to find P(Y=n)

OpenStudy (anonymous):

In that case, my initial thought is: \[ E(X|Y=y)=\sum_{k=1}^{\infty}k\cdot p(k,y) \]

OpenStudy (agent47):

\[P(X=k|Y=n)=\frac{C2^{-k}}{n} \space when \space n \le k \space and \space 0 \space otherwise\]right?

OpenStudy (agent47):

So do we split these into two sums? k=1 to n and k=n to inf?

OpenStudy (anonymous):

You can start \(k\) at \(y\):\[ E(X|Y=y)=\sum_{k=y}^{\infty}k\cdot p(k,y) \]

OpenStudy (anonymous):

This insures that \(n\) is never larger.

OpenStudy (anonymous):

good luck, i gotta go sorry i can't be more help here

OpenStudy (agent47):

no problem satellite, thanks for trying!

OpenStudy (agent47):

@wio so is that sum just:\[\sum_{k=y}^{\infty}{k*\frac{C*2^{-k}}{n}}=\frac{C}{n}\sum_{k=y}^{\infty}{k*2^{-k}}\]

OpenStudy (anonymous):

Well, the \(n=y\).

OpenStudy (agent47):

Oh yea you're right. Also my professor mentioned that the constant C cancels out when you compute conditional expectation.

OpenStudy (anonymous):

My reasoning was this: \[ \Pr(X=k) = p(k, y) \]Then I used expected value definition on that.

OpenStudy (agent47):

Ah I see, So anyway we're left with\[\frac{C}{y}\sum_{k=y}^{\infty}\frac{k}{2^k}\].. Sums are not my strongest suit..

OpenStudy (agent47):

Is that it?

OpenStudy (anonymous):

I think so, it's a rather obscure sum.

OpenStudy (agent47):

So what we have is:\[E(X|Y=y)=\sum_{k=y}^{\infty}k\cdot p(k,y)=\sum_{k=y}^{\infty}{k \cdot \frac{C \cdot 2^{-k}}{y}}=\]\[=\frac{C}{y}\sum_{k=y}^{\infty}{\frac{k}{2^k}}\] which equals to whatever wolframalpha came up with

OpenStudy (anonymous):

Yeah, I think that looks good to me.

OpenStudy (agent47):

Cool, thanks wio. Also dividing that sum by y results in a rather nice-ish result on wolframalpha:\[\frac{2^{1-y}(y+1)}{y}\] http://www.wolframalpha.com/input/?i=(sum(k%2F2%5Ek,+k,+y,+infinity))%2Fy

OpenStudy (agent47):

So the end result ends up being C times that ^

OpenStudy (anonymous):

It's be great if there were a way to test the solution though.

OpenStudy (agent47):

Yea, unfortunately I dont have the answer key :/